OpenStudy (anonymous):

what is the point of having unit vectors? why can't we care just about vector itself, why to add some additional entities?

OpenStudy (baru):

vectors are quantities that have magnitude as well as direction. unit vectors characterize the 'direction' aspect of the vector.

OpenStudy (baru):

for example, consider \(\vec{A} =<3,4>\\\vec{B} =<6,8>\\\vec{C} =<45,60>\)

OpenStudy (baru):

these are three vectors that have different magnitudes, but have the same direction. how can we express this fact as a neat mathematical expression?

OpenStudy (baru):

we use unit vectors, let\[\hat{m}=\frac{<3~,~4>}{\sqrt{3^2+4^2}}=<\frac{3}{5}~,~\frac{4}{5}>\]

OpenStudy (baru):

thus using \(\hat{m}\) we can rewrite our vectors as \[\vec{A}=5 \hat{m} \\ \vec{B}= 10 \hat{m} \\ \vec{C}=75 \hat{m} \]

OpenStudy (baru):

so its not that we have put in any 'additional' entities, its just that we have re-written it in an alternate form by separating its 'magnitude' from its 'direction' all those three vectors have magnitude 5,10 and 75 respectively, and point in the direction \(\hat{m}\).

OpenStudy (phi):

unit vectors are useful if you want the "projection" of some vector in a particular direction u for example A dot u = |A| |u| cos theta when |u|= 1, A dot u gives us the projection also, if we are interested in the angle between two directions, it's more convenient if the directions are represented as unit vectors. then u dot v = |u| |v| cos theta = cos theta or theta = acos ( u dot v)